OpenStudy (anonymous):

Use functions f(x) and g(x) to answer the question. f (x) = √5-x ; g(x) = √x+5 What is the domain of (f/g) (x) Option A: {x|x <5} B: {x | x ≤5} C: {x | -5 ≤ x <5} D: {x | -5 < x ≤ 5}

OpenStudy (anonymous):

im quesing it's option D

OpenStudy (anonymous):

domain of \(f\) is \(5-x\geq 0 \iff x\leq 5\) domain of \(g\) is \(x+5\geq 0\iff x\geq 5\) but since \(g\) is in the denominator it cannot give zero so ...

OpenStudy (anonymous):

yeah D is a good choice

OpenStudy (anonymous):

Thank you